# Electron-ion recombination in composite interactions in liquid xenon

### v1

@@ I Introduction
Liquid xenon is widely used in rare event searches because of its low intrinsic radioactivity, the production of both scintillation and ionization signals, and its ability to identify the types of particle interactions. Currently there are 3 multi-tonne dual-phase xenon Time Projection Chamber (TPC) experiments in operation that search for direct dark matter interactions Aalbers et al. (2024a); Aprile et al. (2024); Bo et al. (2024). A tonne-scale xenon detector with a 136Xe-enriched target has also been proposed to study the hypothesized neutrinoless double beta decay process Adhikari et al. (2021).
@@
The response of liquid xenon to different particle interactions has been extensively studied experimentally, and a comparable effort is made to model these interactions Dahl (2009); Szydagis et al. (2025). In particular, the charge and light outputs of liquid xenon from nuclear recoil (NR) interactions, which represent possible dark matter signals, have been calibrated down to the sub-keV region Lenardo et al. (2019); Akerib et al. (2025) and up to hundreds of keV Pershing et al. (2022), and its responses to electron recoils (ERs) have also been studied with a variety of sources including 37Ar, 3H, and 83mKr XENON Collaboration (2023); Akerib et al. (2016); Boulton et al. (2017); Akerib et al. (2017a). These calibrations have enabled the precise modeling of expected dark matter-induced NR signals in liquid xenon and their strong separation from ambient ER backgrounds. The precision of measurement data and model predictions continues to be improved with ongoing efforts.
@@
Recent progress in xenon-based experiments has drawn attention to a new class of particle interactions, which produces multiple ionizing particles at the same time and the same location. We refer to these events as composite interactions. Different from a “multi-site” interaction that contains two or more (nearly) simultaneous energy depositions at separate locations, particle interactions in a composite event are so close to each other that the production of scintillation and ionization signals by different particles is intertwined. Therefore, the response of xenon to a composite event is not a sum of those for the constituent interactions.
@@
An example of composite interaction is the electron capture (EC) decay, in which a nucleus captures an inner shell electron and emits a neutrino. Because neutrinos usually escape a detector without interacting, the main signature of an EC decay is the atomic relaxation of the shell vacancy and, sometimes, deexcitation gamma rays from the newly formed nucleus. EC decays of 125Xe, 127Xe and 37Ar are widely used in the calibration of xenon detectors for their well-defined, low-energy signatures Akerib et al. (2017b); Temples et al. (2021); Aprile et al. (2017); Akimov et al. (2014); XENON Collaboration (2023); Boulton et al. (2017); Aalbers et al. (2025). However, EC decays usually lead to multiple decay products, making the xenon response complex. For example, the primary decay channel of an L-shell xenon vacancy (or a higher shell) is Auger electron emission. Each Auger decay results in the emission of an electron and the production of two outer shell vacancies, which would also decay primarily through the Auger process. Therefore, a single xenon L-shell EC may produce several low-energy electrons at nearly the same time. The slowdown of each primary particle in liquid xenon produces a group of ions and electrons, some of which may recombine and reduce the measurable ionization signal strength. In an EC decay, the exposure of ions from one primary particle to electrons from other primary particles leads to additional electron-ion recombination. As a result, the observed light and charge outputs of an EC decay can deviate from that of a single electron of the same energy. This effect was observed in XELDA and LZ, with the L-shell EC decays measured to have $\sim$90of the beta charge yield Temples et al. (2021); Aalbers et al. (2025).
@@
The charge suppression effect from recombination enhancement is more significant in events with more complex compositions. Double EC (DEC) decays of 124Xe have been observed in the XENON1T and LZ experiments XENON Collaboration (2019); Aalbers et al. (2024b). 124Xe is a naturally occurring xenon isotope with $\sim$0.1abundance. It can decay to 124Te by capturing two inner shell electrons and emitting two neutrinos, with an extraordinarily long half-life of $>$1022 years. Due to the two simultaneous EC decays at the same location, the light and charge generation process in 124Xe DEC events is strongly modified. In LZ, the observed charge yield of double-L EC decays is only $\sim$64of the sum of two single-L EC decays, or $\sim$70of that of beta events of the same total energy. The reduction in charge yield and increase of light yield cause these events to leak into the NR parameter space more than typical ERs, making them a dangerous background in dark matter searches Aalbers et al. (2024a). Similarly, elastic neutrino scattering with electrons in xenon can produce a recoiling electron and a simultaneous atomic vacancy, another composite event topology with an increased leakage into the dark matter parameter space. As liquid xenon experiments continue to increase exposure and grow in size XLZD Collaboration et al. (2024), the rare 124Xe DEC decays and neutrino-electron scatters will become major backgrounds, and an independent evaluation of the xenon response is required.
@@
Another important signal with a composite signature is the Migdal interaction. The Migdal effect predicts that when a neutral particle scatters off a nucleus, it could result in an atomic excitation or ionization due to the electromagnetic coupling between the nucleus and shell electrons Migdal (1941); Ibe et al. (2018). This signal contains both a NR and an ER component, and the latter further consists of a Migdal electron (in the case of Migdal ionization) and an atomic shell vacancy. The inelastic nature of this process allows a bigger fraction of the neutral particle’s kinetic energy to be transferred, which can make a detector more sensitive to low-energy particles through scattering. Despite the extremely small probability for Migdal interactions Ibe et al. (2018), several experiments have claimed competitive low-mass dark matter sensitivities by assuming this signal channel Akerib et al. (2019); Albakry et al. (2023); Aprile et al. (2019); DarkSide-50 Collaboration et al. (2022). Unfortunately, the effect has not been experimentally verified to date. In addition to other challenges, the composite nature of the Migdal signal creates large uncertainties in the signal response Xu et al. (2024); Bang (2024); Vaitkus (2024), making a definitive search difficult. Even if the Migdal effect is confirmed, dark matter searches with considerable NR energy deposits are still subject to the aforementioned uncertainty in the signal model DarkSide-50 Collaboration et al. (2022).
@@
Direct measurements of liquid xenon’s response to composite interactions are highly desired. However, the only measurements demonstrated so far are the simple single-shell EC decays by 125Xe, 127Xe, and 37Ar Aalbers et al. (2023); Temples et al. (2021); XENON Collaboration (2023); Boulton et al. (2017); Akimov et al. (2014); Aalbers et al. (2025). Calibration of other composite tracks is less practical due to the following reasons. For 124Xe DEC decays, its low isotopic abundance and extraordinarily long half life make a high precision measurement prohibitive; so far, the highest statistics for this process come from LZ, an experiment that suffers from a degeneracy between this background and high-energy signals in some dark matter models Fitzpatrick et al. (2013). Similarly, despite its rarity, a Migdal interaction is by far the most likely mechanism to produce overlapping NRs and ERs. On the other hand, proxy calibrations face substantial difficulty in producing prescribed ionizing particles that perfectly overlap. Compton scatter and photoabsorption of gamma-rays can liberate an electron and simultaneously produce an atomic vacancy, but the probability for inner-shell ionization is low and the signals cannot be conveniently separated from backgrounds. It should be noted that regular “single-track” particle interactions may contain composite tracks when an energetic secondary particle is produced, but this process is stochastic and cannot be controlled. In fact, it is one of the main sources of recombination fluctuation Dahl (2009), which is captured statistically in the measured spreads of charge and light signals.
@@
This work aims to provide a model that evaluates the liquid xenon response to low-energy composite events based on its measured response to the constituent interactions. We take a data-driven approach that has a minimal dependence on xenon microphysics. The model predictions will be compared with available experimental data from composite interactions. This paper is organized as follows. Sec. II describes the method used for the modeling; Sec. III presents the results of the model and their comparison with data; Sec. IV discusses the validity of the assumptions and implementations of the modeling method, and Sec. V concludes this work.
@@ II Modeling Method
A composite event contains multiple ionizing particles emitted from the same location at the same time. Each primary particle produces a positive ion track and a negative electron cloud as it slows down in liquid xenon. In this work, we use the term “track” to refer to a group of ions affiliated with a primary ionizing particle. The movement of positive ions under electric fields or from diffusion is relatively small, so they are treated as stationary. The electrons that carry significant initial kinetic energy will have the energy dissipated through scattering with neighboring xenon atoms and thermalize  [^1] In an external electric field electrons are accelerated by the field while they lose energy through collision with atoms, so the “thermal” energy distribution depends on the electric field.. If an electron wanders close to a positive ion and its kinetic energy does not allow it to escape, it may recombine with the ion. A good understanding of the recombination process for single particle tracks in liquid xenon has been achieved by combining experimental measurements and theoretical modeling Thomas and Imel (1987); Dahl (2009); Szydagis et al. (2025).
We extend the knowledge of how an ion track recombines with its “native” electron cloud to its interactions with “foreign” electrons from a co-located ion track. This approach is motivated by the estimation that ionization electrons can travel 4–5 $\mu$m in liquid xenon before they are thermalized Mozumder (1995), a distance larger than the typical track size of ERs of up to 50 keV or that of NRs up to 100s of keV Dahl (2009). Therefore, the distribution of thermalized electrons with different origins should appear nearly identical to the ion tracks. Electron-ion recombination occurring prior to thermalization is indistinguishable from direct excitations Dahl (2009); the effect of varying exciton-to-ion ratios is a source of systematic uncertainty and will be discussed in Sec. III.
This work extracts the recombination property of an ion track and applies it to the interaction with all available electrons in a composite event. The probability for an ion track to recombine with electrons has a complex dependence on the track geometry, the ionization density, its relative location within the electron cloud and also external electric fields. The instant recombination probability for any individual electron and ion is hard to estimate, but the average recombination probability of all electrons and ions in the track can be empirically calculated from the observed xenon response. For a particle track that, by itself, produces $N_{e}$ electrons and $N_{ph}$ photons, its initial ion number before recombination ($N_{i}$) can be calculated using the exciton-to-ion ratio ($N_{ex}/N_{i}$).
@@
A simple cross-track recombination model can be constructed by assuming that $\bar{r}$ is a constant property of an ion track, independent of the environment that the track is in, so the same $\bar{r}$ value can be applied to both in-track and cross-track electron-ion recombination in composite events. In the case of $m$ co-located ion tracks in a composite event, all $\bar{r}_{k}$} values, where $k=1...m$, can be independently calculated using Eq. 2 if single-track yield data is available. However, these $m$ probabilities are not independent from each other because one electron cannot combine with multiple ion tracks. We therefore interpret $\bar{r}_{k}$ as the conditional probability for an electron to combine with track $k$ when the electron does not combine with any other track. This leads to a renormalization of the recombination probability as follows
@@
Here $P_{k}$ is the normalized probability for an electron to recombine with track $k$. With all {$P_{k}$} values obtained, the multi-track electron-ion recombination process can be modeled as a multinomial distribution, and the number of electrons surviving recombination, or the charge yield of the composite event, can be calculated.
This recombination method is philosophically simple and easy to implement, but can overestimate the electron-ion recombination strength. It is due to the oversimplified constant-$r$ assumption. The ability of an ion track to capture electrons depends on the number of available ions in the track so the average probability $\bar{r}$ is a function of both the initial track ion number and the final ion number. In a composite event, the additional cross-track recombination further reduces the final ion count in a track, which also reduces the average recombination probability $\bar{r}$. Therefore, results from this constant-$r$ method should be seen as a lower bound of the charge yield for a composite interaction.
This weakness of the model can be mitigated by incorporating an electron recombination probability that varies with the track ion number. This approach will be referred to as the running-$r$ method. For single-track events, Eq. 1 and 2 calculate the initial ion count, $N_{i}$, and the average recombination probability, $\bar{r}$, using the observed light and charge yields. In the same way, an empirical function of $\bar{r}(N_{i})$ can be derived with experimental data from single-track events of different energies. This approach assumes that the energy dependence of recombination is driven by the ion number in the event, which is approximately true for a modest change of event energy. An alternative way to get $\bar{r}(N_{i})$ is to use the box recombination model developed by Thomas and Imel (TIB) Thomas and Imel (1987). TIB predicts an analytical formulation for $\bar{r}$ as:
@@
where $\xi$ is an intermediate parameter, $\alpha$ is the recombination coefficient, $a$ is a dimensional parameter, $\mu_{-}$ is the electron mobility, and $E$ is the strength of an external electric field. This model has been used to describe the recombination process in low-energy NR and ER events in liquid xenon Dahl (2009); Szydagis et al. (2025). Both the empirical and TIB recombination functions are studied in this work.
@@
The scaling parameter $\mathcal{C}_{t}$ is needed because $\bar{r}(N_{i})$ is the average recombination probability for all electrons over a range of ion count in the track, which differs from the instant $r(N_{i})$ value at the beginning of the process. Thanks to the monotonic nature of $\bar{r}(N_{i})$, $\mathcal{C}_{t}$ allows $r(N_{i})$ to take a value greater than $\bar{r}(N_{i})$, and as the number of ions decreases $r$ is gradually reduced to below the average recombination probability. In the TIB case, $\mathcal{C}_{t}$ also absorbs the track property parameter $\alpha/(4a^{2}\mu_{-}E)$. The value of $\mathcal{C}_{t}$ for a track can be determined from experimental data through Monte Carlo simulations of the recombination process by testing a range of $\mathcal{C}_{t}$ values. Each $\mathcal{C}_{t}$ corresponds to a unique running-$r$ function and produces a distribution of the surviving electron number for the single-track event of interest; the $\mathcal{C}_{t}$ value producing the best match with the experiment charge yield is chosen and applied to this track in the modeling of composite interactions. When $r(N_{i})$ is applied to the tracks in a component event, the probability renormalization and recombination simulation follow the same approach as used in the constant-$r$ method.

### v2

@@ I Introduction
Liquid xenon is widely used in rare event searches because of its low intrinsic radioactivity, the production of both scintillation and ionization signals, and its ability to identify the types of particle interactions from the energy partition between these two signal channels Aprile et al. (2006). Currently there are 3 multi-tonne dual-phase xenon Time Projection Chamber (TPC) experiments in operation that search for direct dark matter interactions Aalbers et al. (2024a); Aprile et al. (2024); Bo et al. (2024). A tonne-scale xenon detector with a 136Xe-enriched target has also been proposed to study the hypothesized neutrinoless double beta decay process Adhikari et al. (2021).
@@
The response of liquid xenon to different particle interactions has been extensively studied experimentally, and a comparable effort is made to model these interactions Dahl (2009); Szydagis et al. (2025). In particular, the charge and light outputs of liquid xenon from nuclear recoil (NR) interactions, which represent possible dark matter signals, have been calibrated down to the sub-keV region Lenardo et al. (2019); Akerib et al. (2025) and up to hundreds of keV Pershing et al. (2022), and its responses to electron recoils (ERs) have also been studied with a variety of sources including 37Ar, 3H, and 83mKr XENON Collaboration (2023); Akerib et al. (2016); Boulton et al. (2017); Akerib et al. (2017a). These calibrations have enabled the precise modeling of expected dark-matter-induced NR signals in liquid xenon and their strong separation from ambient ER backgrounds. The precision of measurement data and model predictions continues to be improved with ongoing efforts.
@@
Recent progress in xenon-based experiments has drawn attention to a new class of particle interactions, which produces multiple ionizing particles at the same time and the same location. Due to the close proximity, the electrons and ions produced by different ionizing particles overlap and can interact; as a result, the observable light and charge signals by a particle are altered by neighboring particles. We refer to these events as composite interactions. Different from a “multisite” interaction that contains two or more (nearly) simultaneous energy depositions at well-separated locations, a composite event produces a xenon response that is not the sum of those for the constituent interactions.
@@
An example of composite interaction is the electron capture (EC) decay, in which a nucleus captures an inner shell electron and emits a neutrino. Because neutrinos usually escape a detector without interacting, the main signature of an EC decay is the atomic relaxation of the shell vacancy and, sometimes, deexcitation gamma rays from the newly formed nucleus. EC decays of 125Xe, 127Xe and 37Ar are widely used in the calibration of xenon detectors for their well-defined, low-energy signatures Akerib et al. (2017b); Temples et al. (2021); Aprile et al. (2017); Akimov et al. (2014); XENON Collaboration (2023); Boulton et al. (2017); Aalbers et al. (2025). However, EC decays usually lead to multiple decay products, making the xenon response complex. For example, the primary decay channel of an L-shell xenon vacancy (or a higher shell) is Auger electron emission. Each Auger decay results in the emission of an electron and the production of two outer shell vacancies, which would also decay primarily through the Auger process. Therefore, a single xenon L-shell EC may produce several low-energy electrons at nearly the same time. The slowdown of each primary particle in liquid xenon produces a group of xenon ions and electrons, the recombination of which determines the measurable light and charge signal strengths. In an EC decay, the exposure of xenon ions from one primary particle to electrons from other primary particles leads to additional electron-ion recombination. As a result, the signal output of an EC decay can differ from that of a single electron of the same energy. This effect was observed in XELDA and LZ, with the L-shell EC decays measured to have $\sim$90of the beta charge yield Temples et al. (2021); Aalbers et al. (2025).
@@
The charge suppression effect from recombination enhancement is more significant in events with more complex compositions. Double EC (DEC) decays of 124Xe have been observed in the large xenon TPC experiments Aprile et al. (2019a); Aalbers et al. (2024b); Aprile et al. (2022); Bo et al. (2025). 124Xe is a naturally occurring xenon isotope with $\sim$0.1abundance. It can decay to 124Te by capturing two inner shell electrons and emitting two neutrinos, with an extraordinarily long half-life of $>$1022 years. Due to the two simultaneous EC decays at the same location, the light and charge generation process in 124Xe DEC events is strongly modified. In the 2024 WIMP search of LZ (WS2024), the observed charge yield of double-L EC decays is only $\sim$64of the sum of two single-L EC decays, or $\sim$70of that of beta events of the same total energy Aalbers et al. (2024a). The reduction in charge yield and increase of light yield cause these events to leak into the NR parameter space more than typical ERs, making them a dangerous background in dark matter searches. Similarly, elastic neutrino scattering with electrons in xenon can produce a recoiling electron and a simultaneous atomic vacancy, another composite event topology with an increased leakage into the dark matter parameter space. As liquid xenon experiments continue to increase exposure and grow in size Aalbers et al. (2024c), the rare 124Xe DEC decays and neutrino-electron scatters will become major backgrounds, and an independent evaluation of the xenon response is required.
@@
Another important signal with a composite signature is the Migdal interaction. The Migdal effect predicts that when a neutral particle scatters off a nucleus, it could result in an atomic excitation or ionization due to the electromagnetic coupling between the nucleus and shell electrons Migdal (1941); Ibe et al. (2018). This signal contains both an NR and an ER component, and the latter further consists of a Migdal electron (in the case of Migdal ionization) and an atomic shell vacancy. The inelastic nature of this process allows a bigger fraction of the neutral particle’s kinetic energy to be transferred, which can make a detector more sensitive to low-energy particles through scattering. Despite the extremely small probability for Migdal interactions Ibe et al. (2018), several experiments have claimed competitive low-mass dark matter sensitivities by assuming this signal channel Akerib et al. (2019); Albakry et al. (2023); Aprile et al. (2019b); Agnes et al. (2022). Unfortunately, the effect has not been experimentally verified to date. In addition to other challenges, the composite nature of the Migdal signal creates large uncertainties in the signal response Xu et al. (2024); Bang (2024); Vaitkus (2025), making a definitive search difficult. Even if the Migdal effect is confirmed, dark matter searches with considerable NR energy deposits are still subject to the aforementioned uncertainty in the signal model Agnes et al. (2022).
@@
Direct measurements of liquid xenon’s response to composite interactions are highly desired. However, the only measurements demonstrated so far are the simple single-shell EC decays by 125Xe, 127Xe, and 37Ar Aalbers et al. (2023); Temples et al. (2021); XENON Collaboration (2023); Boulton et al. (2017); Akimov et al. (2014); Aalbers et al. (2025). Calibration of other composite tracks is less practical due to the following reasons. For 124Xe DEC decays, its low isotopic abundance and extraordinarily long half-life make a high-precision measurement prohibitive; so far, the highest statistics for this process come from LZ, an experiment that suffers from a degeneracy between this background and high-energy signals in some dark matter models Fitzpatrick et al. (2013). Similarly, despite its rarity, a Migdal interaction is by far the most likely mechanism to produce overlapping NRs and ERs. On the other hand, proxy calibrations face substantial difficulty in producing prescribed ionizing particles that perfectly overlap. Compton scatter and photoabsorption of gamma-rays can liberate an electron and simultaneously produce an atomic vacancy, but the probability for innershell ionization is low and the signals cannot be conveniently separated from backgrounds. It should be noted that regular “single-track” particle interactions may contain composite tracks when an energetic secondary particle is produced, but this process is stochastic and cannot be controlled. In fact, it is one of the main sources of recombination fluctuation Dahl (2009), which is captured statistically in the measured spreads of charge and light signals.
@@
This work aims to provide a model that evaluates the liquid xenon response to low-energy composite events based on its measured response to the constituent interactions. We take a data-driven approach that has a minimal dependence on xenon microphysics. The model predictions will be compared with available experimental data from composite interactions. This paper is organized as follows: Section II describes the method used for the modeling; Sec. III presents the results of the model and their comparison with data; Sec. IV discusses the validity of the assumptions and implementations of the modeling method, and Sec. V concludes this work.
@@ II Modeling Method
A composite event contains multiple ionizing particles emitted from the same location at the same time. Each primary particle produces a positive ion track and a negative electron cloud as it slows down in liquid xenon. In this work, we use the phrase “ion track”, or “track” in short, to refer to a group of ions affiliated with a primary ionizing particle. The movement of positive ions under electric fields or from diffusion is relatively small, so they are treated as stationary. The electrons that carry significant initial kinetic energy will have the energy dissipated through scattering with neighboring xenon atoms and thermalize  [^1] In an external electric field electrons are accelerated by the field while they lose energy through collision with atoms, so the “thermal” energy distribution depends on the electric field strength.. If an electron wanders close to a positive ion and its kinetic energy does not allow it to escape, it may recombine with the ion. A good understanding of the recombination process for single particle tracks in liquid xenon has been achieved by combining experimental measurements and theoretical modeling Thomas and Imel (1987); Dahl (2009); Szydagis et al. (2025).
We extend the knowledge of how an ion track recombines with its “native” electron cloud to its interactions with “foreign” electrons from a colocated ion track. This approach is motivated by the estimation that ionization electrons can travel 4–5 $\mu$m in liquid xenon before they are thermalized Mozumder (1995), a distance larger than the typical track size of ERs of up to 50 keV or that of NRs up to 100s of keV Dahl (2009). Therefore, the distribution of thermalized electrons with different origins should appear nearly identical to the ion tracks. Electron-ion recombination occurring prior to thermalization is indistinguishable from direct excitations Dahl (2009); the effect of varying exciton-to-ion ratios is a source of systematic uncertainty and will be discussed in Sec. III.
This work extracts the recombination property of an ion track and applies it to the interaction with all available electrons in a composite event. This treatment assumes that the spatial separation between the tracks, as well as the individual track length, is significantly shorter than the electron cloud size, beyond which a more sophisticated model will be needed. The probability for an ion track to recombine with electrons has a complex dependence on the track geometry, the ionization density, its relative location within the electron cloud and also external electric fields. The instant recombination probability for any individual electron and ion is hard to estimate, but the average recombination probability of all electrons and ions in the track can be empirically calculated from the observed xenon response. For a particle track that, by itself, produces $N_{e}$ electrons and $N_{ph}$ photons, its initial ion number before recombination ($N_{i}$) can be calculated using the exciton-to-ion ratio ($N_{ex}/N_{i}$).
@@
A simple cross-track recombination model can be constructed by assuming that $\bar{r}$ is a constant property of an ion track, independent of the environment that the track is in, so the same $\bar{r}$ value can be applied to both in-track and cross-track electron-ion recombination in composite events. In the case of $m$ colocated ion tracks in a composite event, all $\bar{r}_{k}$} values, where the track index $k=1...m$, can be independently calculated using Eq. 2 if single-track yield data are available. However, these $m$ probabilities are not independent of each other because one electron cannot combine with multiple ion tracks. We therefore interpret $\bar{r}_{k}$ as the conditional probability for an electron to combine with track $k$ when the electron does not combine with any other track. This leads to a renormalization of the recombination probability as follows
@@
Here $P_{k}$ is the normalized probability for an electron to recombine with track $k$, and the intermediate parameter $A$ provides the normalization. With all {$P_{k}$} values obtained, the multitrack electron-ion recombination process can be modeled as a multinomial distribution, and the number of electrons surviving recombination, or the charge yield of the composite event, can be calculated. Because this method models a composite event based on individual tracks under the same condition, the effects of track topologies and external electric fields are automatically accounted for.
This recombination method is simple to understand and easy to implement, but can overestimate the electron-ion recombination strength due to the oversimplified constant-$r$ assumption. The ability of an ion track to capture electrons depends on the number of available ions in the track so the average probability $\bar{r}$ is a function of both the initial track ion number and the final ion number. In a composite event, the additional cross-track recombination further reduces the final ion count in a track, which also reduces the average recombination probability $\bar{r}$. Therefore, results from this constant-$r$ method should be seen as a lower bound of the charge yield for a composite interaction.
This weakness of the model can be mitigated by incorporating an electron recombination probability that varies with the instant track ion number. This approach will be referred to as the running-$r$ method. For single-track events, Eq. 1 and 2 calculate the initial ion count, $N_{i}$, and the average recombination probability, $\bar{r}$, using the observed light ($N_{ph}$) and charge yields ($N_{e}$). In the same way, an empirical function of $\bar{r}(N_{i})$ can be derived with experimental data from single-track events of different energies. This approach assumes that the energy dependence of recombination is driven by the ion number in the event, which is approximately true for a modest change of event energy. An alternative way to get $\bar{r}(N_{i})$ is to use the box recombination model developed by Thomas and Imel (TIB) Thomas and Imel (1987). TIB predicts an analytical formulation for $\bar{r}$ as:
@@
where $\xi$ is an intermediate parameter, $\alpha$ is the recombination coefficient (recombination probability per ion per unit time), $a$ is a dimensional parameter, $\mu_{-}$ is the electron mobility, and $E$ is the strength of an external electric field. This model has been used to describe the recombination process in low-energy NR and ER events in liquid xenon Dahl (2009); Szydagis et al. (2025). Both the empirical and TIB recombination functions are studied in this work.
Figure 1: Illustration of the running-$r$ recombination modeling process. The dashed box highlights actions on each constituent of a composite track: $N_{i}$ is calculated from track yields; $\mathcal{C}_{t}$ is obtained by testing a range of $C_{t}$ using the 1-track modeling process and finding the value that produces the correct track yield. The constant-$r$ process differs in the red boxes: instead of $\mathcal{C}_{t}$, the average $\bar{r}$ (Eq. 2) is used in each subsequent loop until $N_{i}$ decreases to 0 (then $r=$0).
@@
The scaling parameter $\mathcal{C}_{t}$ is needed because $\bar{r}(N_{i})$ is the average recombination probability for all electrons over a range of ion counts in the track, which differs from the instant $r(N_{i})$ value at the beginning of the process. Thanks to the monotonic nature of $\bar{r}(N_{i})$, $\mathcal{C}_{t}$ allows $r(N_{i})$ to take a value greater than $\bar{r}(N_{i})$, and as the number of ions decreases $r$ is gradually reduced to below the average recombination probability. In the TIB case, $\mathcal{C}_{t}$ also absorbs the track property parameter $\alpha/(4a^{2}\mu_{-}E)$. The value of $\mathcal{C}_{t}$ for a track can be determined from experimental data through Monte Carlo simulations of the recombination process by testing a range of $\mathcal{C}_{t}$ values. Each $\mathcal{C}_{t}$ corresponds to a unique running-$r$ function and produces a distribution of the surviving electron number for the single-track event of interest; the $\mathcal{C}_{t}$ value producing a mean charge yield that matches experimental data is chosen and applied to this track in the modeling of composite interactions. When $r(N_{i})$ is applied to the tracks in a composite event, the probability renormalization and recombination simulation follow the same approach as used in the constant-$r$ method. This process is illustrated in Fig. 1.